# Observing Double White Dwarfs with the Lunar GW Antenna

### v1

@@ 3 Generation of the synthetic population
In order to simulate with accuracy the response of LGWA to the DWD population, inside and outside our Galaxy, it is prominent to generate a reliable synthetic population. The adopted procedure is explained in this section.
@@ 3.1 Population size estimation
In order to generate a realistic population, it is important to estimate the number of DWD systems in our galaxy and in the local Universe. Our analysis assumes that every super-Chandrasekhar DWD system leads to a SN Ia, and therefore the SN Ia rate derived from this assumption should be considered a lower limit, since other final outcomes are possible from a super-Chandrasekhar DWD merging (Luo et al., 2025). However, two opposite effects must be considered too: firstly, some SNe Ia may originate from sub-Chandrasekhar mass binaries (Badenes & Maoz, 2012; Ruiter & Seitenzahl, 2025; Scalzo et al., 2014), and secondly, some of the SNe Ia may result from the SD scenario, in a mixed progenitor scenario (Ruiter & Seitenzahl, 2025). At this stage, it is not realistically possible to determine how these different contributions balance out. GW observations will also contribute to clarifying this aspect.
The adopted SN Ia rate in our Galaxy is $r=(5.4\pm 1.2)\cdot 10^{-3}$ yr-1 (Li et al., 2011), and is considered constant over time: this approximation is valid on the timescales of the detectable DWDs (those characterized by a GW frequency above 1 mHz) because the time required to merge starting from $f_{i}=1$ mHz is around 1 – 3 Myr depending on the masses, sufficiently small compared with the entire Hubble time (13500 Myr) to consider the SN Ia rate locally fixed.
The density distribution in frequency space of the DWD population required to sustain this rate is obtained in the following.
@@
which integrated in a frequency interval $I_{f}=(f_{m},f_{M})$ gives
$N(I_{f})=\frac{r}{\Delta(\mathcal{M})}\bigl{(}f_{m}^{-\frac{8}{3}}-f_{M}^{-\frac{8}{3}}\bigr{)}$ (9)
This formula provides the expected number of galactic DWDs which are currently spiralling with a frequency between $f_{m}$ and $f_{M}$. After eq. 9 and the adopted SN Ia rate, we infer a total number of DWD equal to $(2.0\pm 0.5)\cdot 10^{4}$ for the frequency band between 1 mHz and 5 mHz, $228\pm 62$ DWDs between 5 mHz and 1 dHz and $42\pm 11$ between 1 dHz and 5 dHz. The reported errors are calculated propagating both the error on the rate estimation and the statistical dispersion of the function $\Delta(\mathcal{M})$ (defined in eq. 5). This dispersion was obtained from the galactic DWD population obtained in sec. 3.3, considering only the systems with $M_{tot}>M_{Ch}$. As the errors are calculated from a mixed approach, they are intended to be indicative.
@@ 3.2 Primordial population and evolution
Being the DWD parameter distribution poorly known, the present stellar population can be generated evolving a primordial population with the code SeBa [^4] https://github.com/amusecode/SeBa. This code makes it possible to completely simulate the evolution of a stellar population; in particular, for the sake of this analysis, it is used to evolve a population of binaries in the parameter ranges where the formation of a DWD is possible. We sample the population parameters, namely the masses $m_{1}$ and $m_{2}$, the initial orbital separation, the eccentricity and the metallicity $Z$ accordingly with some distributions implemented in the code. The fiducial parameter distributions and parameter intervals are chosen accordingly to Korol et al. (2017) and reported in tab. 1. Despite effectively using a thermal distribution for the eccentricity, all the short-period DWDs show $\varepsilon=0$ at the formation of the DWD, at the end of the CE phase. Thus the effects of eccentricity will be neglected, considering only circular orbits, as anticipated in sec. 2.2. After the sampling, every binary system in the population is evolved independently for a chosen time interval, taking into account all the major effects that contribute to the evolution of the single stars and the system as a whole. The state of the system for very relevant evolutionary phase is stored.
@@
The resulting population consists of $2\cdot 10^{7}$ binary systems, each with its evolutionary history over a 13.5 Gyr evolution time interval. Approximately $75\%$ of the systems become DWD at some point of the evolution; Among these, only $\approx 5.8\%$ at the end of the CE phase display an orbital separation which is sufficiently small in order to enter the mHz frequency band at some time before the end of the 13.5 Gyr evolutionary period, due to the GW emission. Finally, only few hundreds fall in the LGWA frequency band exactly at the end of the evolutionary period, all around the mHz and none above 10 mHz. Thus the evolution of a single burst cannot contain a statistically relevant sample of the population, since the expected DWD galactic population in the mHz should account for $\approx 2\cdot 10^{4}$ systems above $M_{\text{Ch}}$ in order to reproduce the current SN Ia rate (see the estimations in sec. 3.1), and the simulation can’t reproduce at all the rare but crucial population with $f_{GW}>0.01$ Hz in the most sensitive band of LGWA.
The solution comes from using the generated population as a $\delta$ function and convolve it with a chosen star formation history (SFH). For every $t$, $0<t<13.5$ Gyr $\equiv t_{0}$, in the population a certain number of new systems are formed (systems/year); it is thus possible to define a formation rate $R_{\text{SFH}}(t)\in[0,1]$ by normalizing the SFH of the population at the value that it reached at the peak of star formation. For the time interval $(t,t+\delta t)$ a fraction $R_{\text{SFH}}(t)$ of the whole SeBa population is sampled and the frequency is evaluated at the present time, namely at the time $\tilde{t}=t_{0}-t+\xi\cdot\delta t$ in the simulation. The parameter $\xi$ is drawn randomly from the interval (0.5,0.5)(-0.5,0.5) for every system to account for a uniform distribution within the minimum period $\delta t$, and thus smoothing the $\delta$ functions into a continuous distribution. If the algorithm finds a system that for $\tilde{t}$ lies in the LGWA frequency band, the system is added to the convolved present population. In addition, let $M_{s}$ be the total simulation mass, namely the sum of the masses of the stars at the time of formation ($\tilde{t}=0$); this characteristic mass is used in sec. 3.3 to correctly account for the various stellar populations within the Milky Way (MW).  [^5] This is the most reliable estimation of the convolved population mass, even if it does not represent the effective mass at the end of the evolution. Obtaining this second estimation would require the complete simulation of the entire population evolution, which is not reasonable considered the necessary computational cost.
This approach solves the initial problem as not only the few short-period DWDs that remain at $t=13.5$ Gyr are considered, but all the systems that at some point of the 13.5 Gyr evolution had become a DWD. In this way the computationally expensive SeBa simulation is better exploited. In addition, the $\delta$-convolving approach benefits from two main advantages:
However, there are some drawbacks that must be taken into account:
  • The $\delta$ burst refers to fixed initial parameters that could change over time in the real SFH. For example, the metallicity is expected to increase as the galaxy ages or vary for the different galactic populations (bulge, thin disk, thick disk, see sec. 3.5), but this effect cannot be taken into account since a variation in the metallicity would require to re-run the evolutionary simulation. The same is true for the IMF and the binary fraction.
@@ 3.3 Star Formation Histories
Once evolved the $\delta$ population sample, the final population is generated via the convolution with a chosen SFH. The spatial distribution of the DWDs (see sec. 3.5) needs to be performed considering distinct components of the total galactic population. We consider a central bulge, a thin disk and a thick disk as described in sec. 3.5. In principle, also other stellar populations could be simulated, as the Magellanic clouds, the single globular clusters, and finally other galaxies within the horizon of LGWA, to completely model its detection capabilities. However, for this analysis an accurate SFH model is developed for the MW and than reused to model also the extragalactic DWD population.
The SFH of the various components is usually derived from the chemical analysis of the stellar populations; this results in a series of star formation rates at any time from formation up to present days. To simulate the SFH it is chosen a simple analytical approximant that closely resembles the experimental data. The discrepancy between the approximant and the true SFH is negligible, as the global properties of the convolved population are (a posteriori) very robust. The important characteristics of the SFH are the overall abundance and the approximate distribution in time. The SFH of the disks is chosen in accordance with Snaith et al. (2014), Snaith et al. (2015), and presents an early phase of bursty star formation in the thick disk, followed by a clear gap approximately 7.5 Gyr ago, and a less intense phase in which the thin disk is formed. In fig. 2 are shown the original data from Snaith et al. (2014) and the two approximants for the disks superimposed.
The SFH of the bulge is chosen accordingly to Lian et al. 2020 (based on CCSNe yields by Kobayashi et al. 2006) to be a Gaussian burst with $\mu=1$ Gyr, $\sigma=0.65$ Gyr, followed by a low-intensity exponential tail $\text{SFH}_{\text{late}}=0.1\exp{\Big{(}-\frac{t}{8.3\,\text{Gyr}}}\Big{)}$ (normalized with respect to the peak of the Gaussian burst). Note that the tail, being nearly quiescent, is less important from the point of view of the total stellar mass produced in the overall SFH, but it is relevant for the final number of present-day DWDs: the binary systems that formed from stars originated in the burst are more likely to have already coalesced. In fact, the fraction of the DWDs in the bulge that were formed in the exponential tail account for the 40% of the bulge DWD population. A complete consensus on the SFH of the various components is not yet established: for example, regarding the bulge, an opposite model is presented by Haywood et al. (2016), with a more intense SFR in the late phase. In principle, with the straightforward procedure described here it is simple to compare different possible SFHs.
@@
The DWD relative abundance among the components is chosen comparing the mass of the simulations $M_{s,\text{bulge}}$, $M_{s,\text{thin}}$, $M_{s,\text{thick}}$ with the estimated masses reported in tab. 2, therefore there is not the problem of weighting the SFHs of the various components via the multiplicative density coefficients.
The ratio between the population masses is obtained by properly sampling the convolved populations. Let $\chi_{b}$, $\chi_{t}$, $\chi_{T}$ be the three sampling coefficients, and $N_{b}$, $N_{t}$, $N_{T}$ the number of super-Chandrasekhar binaries in the convolved simulations, respectively for the bulge, the thin disk and the thick disk; the sampling coefficient are found by imposing the ratio constrains on the masses:
$\frac{\chi_{t}M_{s,\text{thin}}}{\chi_{T}M_{s,\text{thick}}}=\frac{M_{\text{thin}}}{M_{\text{thick}}}\hskip 56.9055pt\frac{\chi_{t}M_{s,\text{thin}}}{\chi_{b}M_{s,\text{b}}}=\frac{M_{\text{thin}}}{M_{b}}$
and accounting for the total number of super-Chandrasekhar DWD, $\chi_{t}N_{t}+\chi_{T}N_{T}+\chi_{b}N_{b}=N_{tot}=2\cdot 10^{4}$ after the estimations presented in sec. 3.1.
Although the normalization of the overall SFH for the three components has been chosen a posteriori via the total mass and the abundance of super-Chandrasekhar DWDs, the ratio between the sampling coefficients of thin and thick disks matches perfectly (better than $\tcperthousand1\tcperthousand$) the ratio between the peaks of the SFHs as given by Snaith et al. (2014), also reported in fig. 2. This demonstrates the robustness of the procedure and the validity of the chosen SFHs. For the bulge, the direct comparison with the original SFH provided in Lian et al. (2020) is more difficult, since the conversion from surface SFR density and total SFR is dependent on the chosen spatial distribution (whereas the normalized SFH is not, namely the conversion occurs via a simple multiplicative coefficient). However, since also the bulge SFH influences the sampling coefficients of the disks, this is an indication that also the treatment for the bulge is correct.
The mass and frequency distributions of the three populations are plotted and commented in fig. 3.
@@ 3.4 Other parameters
To carry out a complete GW analysis with GWFish, it is necessary to fix additional source parameters. The assumptions adopted are listed below:
  • System inclination: the inclination $\theta$ is sampled with an uniform distribution in $\cos\theta$.
  • Polarization angle: $\psi$ is sampled uniformly in $(0,2\pi)$.
  • Phase: sampled uniformly in $(0,2\pi)$, despite being not important for long-period observations.
  • $a_{1}$ and $a_{2}$: components of the WDs spins along the orbital angular momentum, in units of $M^{2}$, where $M$ is the WD mass, in natural units; sampled uniformly in $(0,0.1)$.
@@ 3.5 Spatial distribution within MW
To perform the elaboration, only $r^{\prime}<3$ kpc for the bulge and $R<35$ kpc for the disks are considered. Despite being distinguishable by the chemical composition, this double disk modeling is used exclusively to account for a better spatial distribution.
@@ 3.6 Extragalactic sources: spatial distribution and DWD populations
While for the MW the populations are distributed following a complex spatial model, for the extragalactic sources this is superfluous: as the distance of a galaxy grows larger than its size, the spatial distribution of the population becomes negligible in determining the detectability and PE of the single DWD. Thus, these sources are associated directly with the position of the respective galaxy centers. For each galaxy within 30 Mpc (the estimated horizon of LGWA as described in Appendix A) we extract the following data from the HyperLeda [^7] http://leda.univ-lyon1.fr/ catalog (Makarov et al., 2014):
@@
  • Morphological type of the galaxy, expressed in the Hubble morphological classification. This parameter is used to obtain the SN Ia rate in the case that the color is not available (in particular when the $K$ band lacks, which occurs for a non-negligible fraction of the population).
The SN Ia rate for each galaxy in our sample is estimated based on observed SN Ia rates in the local Universe. These rates are provided as functions of various galaxy properties, including luminosity in the $B$ and $K$ bands, stellar mass, and morphological type. Specifically, we adopt the rate–size relations described in Li et al. (2011), and we use two parameters: the $B$ band luminosity and one chosen among the morphological type and the BKB-K color. The rates are calculated preferably from the BKB-K color (obtained as $m_{B}-m_{K}$) and the $B$ band luminosity. Where this was not possible due to the lacking of the $K$ apparent magnitude, the morphological type and the $B$ band luminosity are used. The $B$ luminosity is $L_{B}=10^{-0.4(M_{B}-M_{B,\odot})}L_{\odot}$, where $M_{B}$ is the absolute corrected $B$ band magnitude, $M_{B,\odot}=5.31\pm 0.03$ mag is the absolute magnitude of the Sun in the $B$ band and $L_{\odot}$ is the luminosity of the Sun. The $B$ luminosity is than expressed in units of $10^{10}L_{\odot}$ as a reference scale unit for the luminosity of galaxies.
Galaxies lacking both data sets are a minor component (mainly negligible stellar aggregates) and are not expected to significantly affect the overall SN Ia rate outside the Milky Way. Therefore, they are excluded from the analysis.
The rate-size relations that relate the B band luminosity and galaxy color or morphological type to the rate $r$ are found to be power laws:
$r(\text{type},L_{B})=\text{SNuB}(\text{type},L_{0})\cdot L_{B}\cdot\bigg{(}\frac{L_{B}}{L_{B0}}\bigg{)}^{\text{RSS}_{B}}$ (10)
$r(B-K,L_{B})=\text{SNuB}(B-K,L_{0})\cdot L_{B}\cdot\bigg{(}\frac{L_{B}}{L_{B0}}\bigg{)}^{\text{RSS}_{B}}$ (11)
Where SNuB is the rate for galaxy with luminosity $L=10^{10}L_{\odot}$, and RSSB is a power-like correction to the simple linear relation between the size (luminosity) and the rate, gauged with respect to a ”typical galaxy” with luminosity $L_{B0}=2.2\cdot 10^{10}\,L_{\odot}$. The average-size SNuB coefficients and the relative RSSB for the different color and morphological types bins, drawn from Li et al. (2011), are listed  [^9] Note that since in the HyperLeda catalog $H_{0}=70\,\text{km s}^{-1}\text{Mpc}^{-1}$ is assumed, while in Li et al. (2011) $H_{0}=73\,\text{km s}^{-1}\text{Mpc}^{-1}$, the rates have been corrected by a factor $q=\Big{(}\frac{H_{0,\text{HyperLeda}}}{H_{0,\text{Li}}}\Big{)}^{2}\simeq 0.919$ via $\text{SNuB}_{\text{corr}}=\text{SNuB}_{\text{Li}}\cdot q\hskip 19.91684pt\text{and}\hskip 19.91684ptL_{B0,\text{corr}}=L_{B0,\text{Li}}\cdot q^{-1}$ in tab. 3. The rates are expressed in units of one SN Ia every 100 years. A plot of the positions of the galaxies with the corresponding rates and distances is shown in fig. 4.
@@
The estimation of cumulative SN Ia rate within 10 Mpc is $31\pm 7$ SN/100 yrs, $245\pm 67$ SN/100 yrs for a 20 Mpc horizon and $650\pm 157$ SN/100 yrs for a 30 Mpc horizon. The errors are computed via a Monte Carlo error estimation starting from the errors on the parameters given by Li et al. (2011), and a 20% correction has been applied to account for the Zone of Avoidance (see sec. 3.6.1). These results, accounting for the statistical errors, are in discrete accordance with the rates reported in Ajith et al. (2025), based on an independent approach. A comparison between the rates obtained in this paper and the rates reported by Ajith et al. (2025) is presented in fig. 5. The figure presents also a comparison with the rates obtained by using preferably the morphological type instead of the color [^10] The BKB-K color is still used where the morphological type is not available.. The relative discrepancy between the two measurements is shown in gray, and is compatible with 0 within the errors. We observe the rates obtained with the type to be consistently more abundant by 5% for $d_{L}>20$ Mpc.
Figure 5: Comparison between rates obtained in this work (with 20% avoidance zone correction factor (AZC), see sec. 3.6.1) and LGWA White Paper (Ajith et al., 2025). We report the rate calculated preferably from the BKB-K color (used in the rest of the paper, here in red) and the rate obtained using preferably the morphological type (here in green). The black line shows the discrepancy in percent between the two estimations. The error bands are obtained with Monte Carlo sampling using the statistical errors on the parameters given by Li et al. (2011).
@@
Since the detectable population outside our Galaxy consists only in the most massive DWDs that experience the merging (see sec. 4), only super-Chandrasekhar merging systems are considered. The realistic number of mergers expected is not sufficient to have a significant statistics, so a higher number is considered and the results are to be intended as probabilistic, once normalized to the true number of expected events. A broader population is obtained by considering all the merging events produced by the synthetic population in a time span from the nominal observation time $t_{0}$ ($t_{0}=13.5$ Gyr as a cosmic time) to $t_{0}+2.4$ Myr. The time interval $\Delta t=2.4$ Myr is chosen observing that the merger rate of the synthetic population is stable for 2.4 Myr and than drops, due to the fact that the population is depleted from the DWD that merged in the meantime. The merger rate is determined by evolving the separation between the binaries only under GW energy dissipation and considering a Roche scenario as the merging condition. From the resulting population we choose to sample a merging population 500 times larger than the expected one (without incompleteness correction, see sec. 3.6.1).
3.6.1 Error budget on extragalactic population

### v2

@@ 3 Generation of the synthetic population
In order to simulate with accuracy the response of LGWA to the DWD population, inside and outside our Galaxy, it is prominent to generate a reliable synthetic population. The adopted procedure is summarized in this section; further details and explanations are reported in Appendix B.
@@ 3.1 Population size estimation
In order to generate a realistic population, it is important to estimate the number of DWD systems in our Galaxy and in the local Universe. Our analysis assumes that every super-Chandrasekhar DWD system leads to a SN Ia, and therefore the SN Ia rate derived from this assumption should be considered a lower limit, since other final outcomes are possible from a super-Chandrasekhar DWD merging (Luo et al., 2025). However, two opposite effects must be considered too: firstly, some SNe Ia may originate from sub-Chandrasekhar mass binaries (Badenes & Maoz, 2012; Ruiter & Seitenzahl, 2025; Scalzo et al., 2014; Liu et al., 2023), and secondly, some of the SNe Ia may result from the SD scenario, in a mixed progenitor scenario, as well as other proposed scenarios (Ruiter & Seitenzahl, 2025; Liu et al., 2023). At this stage, it is not realistically possible to determine how these different contributions balance out. GW observations will also contribute to clarifying this aspect.
The adopted SN Ia rate in our Galaxy is $r=(5.4\pm 1.2)\cdot 10^{-3}$ yr-1 (Li et al., 2011), and is considered constant over time: this approximation is valid on the timescales of the detectable DWDs (those characterized by a GW frequency above 1 mHz) because the time required to merge starting from $f_{i}=1$ mHz is around 1 to 3 Myr depending on the masses, sufficiently small compared with the entire Hubble time (13.5 Gyr) to consider the SN Ia rate locally fixed. In the following, we derive the density distribution in frequency space of the DWD population necessary to sustain this rate.
@@
We infer a total number of super-Chandrasekhar DWD equal to $(2.0\pm 0.5)\cdot 10^{4}$ for the frequency band between 1 mHz and 5 mHz, $228\pm 62$ DWDs between 5 mHz and 1 dHz and $42\pm 11$ between 1 dHz and 5 dHz. The reported errors are calculated propagating both the error on the rate estimation and the statistical dispersion of the function $\Delta(\mathcal{M})$ (defined in eq. 5). This dispersion is computed from the Galactic DWD population obtained in Sect. 3.3, considering only super-Chandrasekhar systems. As the errors are calculated from a mixed approach, they are intended to be indicative.
@@ 3.2 Primordial population and evolution
Being the DWD parameter distribution poorly known, the present stellar population can be generated evolving a primordial population with the code SeBa [^4] https://github.com/amusecode/SeBa. This code makes it possible to completely simulate the evolution of a stellar population; in particular, for the sake of this analysis, it is used to evolve a population of binaries in the parameter ranges where the formation of a DWD is possible. We sample the population parameters, namely the masses $m_{1}$ and $m_{2}$, the initial orbital separation, the eccentricity and the metallicity $Z$ accordingly with some distributions implemented in the code. The fiducial parameter distributions and parameter intervals are chosen accordingly to Korol et al. (2017) and reported in tab. 1. Despite effectively using a thermal distribution for the eccentricity, all the short-period DWDs show $\varepsilon=0$ at the formation of the DWD, at the end of the second CE phase. Thus the effects of eccentricity will be neglected, considering only circular orbits, as anticipated in Sect. 2.2. The CE treatment is implemented in SeBa following the $\gamma\alpha$-prescription, which is thoroughly motivated and discussed in Toonen et al. (2012). This prescription reproduces well the observations specifically in the context of DD SNIa progenitors, which is the main target population of this paper. After the sampling, every binary system in the population is evolved independently for a chosen time interval, taking into account all the major effects that contribute to the evolution of the single stars and the system as a whole. The state of the system for every relevant evolutionary phase is stored.
@@
The resulting population consists of a sample of binary systems, each with its evolutionary history over a 13.5 Gyr evolution time interval. This corresponds to a $\delta$-burst”, from which we obtain the simulated populations through a convolution with the respective star formation histories (SFH).
For every $t$, $0<t<13.5$ Gyr $\equiv t_{0}$, in the population a certain number of new systems are formed (systems/year); it is thus possible to define a formation rate $R_{\text{SFH}}(t)\in[0,1]$ by normalizing the SFH of the population at the value that it reached at the peak of star formation. For the time interval $(t,t+\delta t)$ a fraction $R_{\text{SFH}}(t)$ of the whole SeBa population is sampled and the frequency is evaluated at the present time, namely at the time $\tilde{t}=t_{0}-t+\xi\cdot\delta t$ in the simulation. The parameter $\xi$ is drawn randomly from the interval (0.5,0.5)(-0.5,0.5) for every system to account for a uniform distribution within the minimum period $\delta t$, and thus smoothing the $\delta$ functions into a continuous distribution. If the algorithm finds a system that for $\tilde{t}$ lies in the LGWA frequency band, the system is added to the convolved present population. In addition, let $M_{s}$ be the “total simulation mass”, namely the sum of the masses of the stars at the time of formation ($\tilde{t}=0$); this characteristic mass is used in Sect. 3.3 to correctly account for the various stellar populations within the Milky Way (MW).  [^5] This is the most reliable estimation of the convolved population mass, even if it does not represent the effective mass at the end of the evolution. Obtaining this second estimation would require the complete simulation of the entire population evolution, which is not reasonable considered the necessary computational cost. Further details and motivations for the use of this procedure are discussed in Appendix B.1.
@@ 3.3 Star Formation Histories
Once evolved the $\delta$ population sample, the final population is generated via the convolution with a chosen SFH. The spatial distribution of the DWDs (see Sect. 3.5) needs to be performed considering distinct components of the total Galactic population. We consider a central bulge, a thin disk and a thick disk as described in Sect. 3.5. In principle, also other stellar populations could be simulated, as the Magellanic clouds, the single globular clusters, and finally other galaxies within the horizon of LGWA, to completely model its detection capabilities. However, for this analysis an accurate SFH model is developed for the MW and than reused to model also the extragalactic DWD population.
The SFH of the various components is usually derived from the chemical analysis of the stellar populations; this results in a series of star formation rates at any time from formation up to present days. To simulate the SFH it is chosen a simple analytical approximant that closely resembles the experimental data. The discrepancy between the approximant and the true SFH is negligible, as the global properties of the convolved population are (a posteriori) very robust. The important characteristics of the SFH are the overall abundance and the approximate distribution in time. The SFH of the disks is chosen in accordance with Snaith et al. (2014, 2015), and presents an early phase of bursty star formation in the thick disk, followed by a clear gap approximately 7.5 Gyr ago, and a less intense phase in which the thin disk is formed. Figure 2 shows the original data from Snaith et al. (2014) with the two approximants for the disks superimposed.
The SFH of the bulge is chosen accordingly to Lian et al. 2020 (based on CCSNe yields by Kobayashi et al. 2006) to be a Gaussian burst with $\mu=1$ Gyr, $\sigma=0.66$ Gyr, followed by a low-intensity exponential tail $\text{SFH}_{\text{late}}=0.1\exp{\Big(-\frac{t}{8.3\,\text{Gyr}}}\Big)$ (normalized with respect to the peak of the Gaussian burst). Note that the tail, being nearly quiescent, is less important from the point of view of the total stellar mass produced in the overall SFH, but it is relevant for the final number of present-day DWDs: the binary systems that formed from stars originated in the burst are more likely to have already coalesced. In fact, the fraction of the DWDs in the bulge that were formed in the exponential tail account for the 40% of the bulge DWD population. A complete consensus on the SFH of the various components is not yet established: for example, regarding the bulge, an opposite model is presented by Haywood et al. (2016), with a more intense SFR in the late phase. In principle, with the straightforward procedure described here it is simple to compare different possible SFHs.
@@
The DWD relative abundance among the components is chosen comparing the mass of the simulations $M_{s,\text{bulge}}$, $M_{s,\text{thin}}$, $M_{s,\text{thick}}$ with the estimated masses reported in tab. 2 and accounting for the total number of super-Chandrasekhar DWD, $N_{tot}=2\cdot 10^{4}$ after the estimations presented in Sect. 3.1.
Although the normalization of the overall SFH for the three components has been chosen a posteriori via the total mass and the abundance of super-Chandrasekhar DWDs, the ratio between the sampling coefficients of thin and thick disks matches perfectly (better than $1\tcperthousand$) the ratio between the peaks of the SFHs as given by Snaith et al. (2014), also reported in fig. 2. This demonstrates the robustness of the procedure and the validity of the chosen SFHs. For the bulge, the direct comparison with the original SFH provided in Lian et al. (2020) is more difficult, since the conversion from surface SFR density and total SFR is dependent on the chosen spatial distribution (whereas the normalized SFH is not, namely the conversion occurs via a simple multiplicative coefficient). However, since also the bulge SFH influences the sampling coefficients of the disks, this is an indication that also the treatment for the bulge is accurate.
The mass and frequency distributions of the three populations are plotted and commented in fig. 10. Note that the differences in abundance and distribution are due only to the chosen SFH, since the three populations are obtained starting from the same original $\delta$ burst.
@@ 3.4 Other parameters
To carry out a complete GW analysis with GWFish it is necessary to fix additional source parameters. The system inclination $\cos\theta$, the polarization angle and the phase are sampled uniformly in the respective domains, the components $a_{1}$ and $a_{2}$ of the WDs spins along the orbital angular momentum in units of $m_{1,2}^{2}$ are sampled uniformly between 0 and 0.1.
@@ 3.5 Spatial distribution within MW
Despite being distinguishable by the chemical composition, this double disk modeling is used exclusively to account for a better spatial distribution.
@@ 3.6 Extragalactic sources: spatial distribution and DWD populations
While for the MW the populations are distributed following a multi-component spatial model, for the extragalactic sources this is superfluous: as the distance of a galaxy grows larger than its size, the spatial distribution of the population becomes negligible in determining the detectability and PE of the single DWD. Thus, these sources are associated directly with the position of the respective galaxy centers. For each galaxy within 30 Mpc (the estimated horizon of LGWA as described in Appendix A) we extract the following data from the HyperLeda [^6] http://leda.univ-lyon1.fr/ catalog (Makarov et al., 2014):
@@
  • Morphological type of the galaxy, expressed in the Hubble morphological classification.
The SN Ia rate for each galaxy in our sample is estimated based on observed SN Ia rates in the local Universe. These rates are provided as functions of various galaxy properties, including luminosity in the $B$ and $K$ bands, stellar mass, and morphological type. Specifically, we adopted the rate–size relations described in Li et al. (2011), and we used two parameters: the $B$ band luminosity and one chosen among the morphological type and the BKB-K color. The rates were calculated preferably from the BKB-K color (obtained as $m_{B}-m_{K}$) and the $B$ band luminosity. Where this was not possible due to the lacking of the $K$ apparent magnitude, which is the case for a non-negligible fraction of the population, the morphological type and the $B$ band luminosity were used. See App. B.3 for further details. The $B$ luminosity is $L_{B}=10^{-0.4(M_{B}-M_{B,\odot})}L_{\odot}$, where $M_{B}$ is the absolute corrected $B$ band magnitude and $M_{B,\odot}=5.31\pm 0.03$ mag is the absolute magnitude of the Sun in the $B$ band. The $B$ luminosity is than expressed in units of $10^{10}L_{\odot}$ as a reference scale unit for the luminosity of galaxies.
Galaxies lacking both data sets are a minor component (mainly negligible stellar aggregates) and are not expected to significantly affect the overall SN Ia rate. Therefore, they are excluded from the analysis.
The rate-size relations that relate the B band luminosity and galaxy color or morphological type to the rate $r$ are phenomenologically parametrized by (Li et al., 2011):
$r(L_{B})=\text{SNuB}(L_{0})\cdot L_{B}\cdot\bigg(\frac{L_{B}}{L_{B0}}\bigg)^{\text{RSS}_{B}}$
Where SNuB is the rate for a galaxy with luminosity $L=10^{10}L_{\odot}$, and RSSB is a power-like correction to the simple linear relation between the size (luminosity) and the rate, gauged with respect to a “typical galaxy” with luminosity $L_{B0}=2.2\cdot 10^{10}\,L_{\odot}$. Both the coefficients depend also on the choice of either color or morphological type as the secondary parameter. The average-size SNuB coefficients and the relative RSSB from Li et al. (2011) are listed  [^8] Note that since in the HyperLeda catalog $H_{0}=70\,\text{km s}^{-1}\text{Mpc}^{-1}$ is assumed, while in Li et al. (2011) $H_{0}=73\,\text{km s}^{-1}\text{Mpc}^{-1}$, the rates have been corrected by a factor $q=\Big(\frac{H_{0,\text{HyperLeda}}}{H_{0,\text{Li}}}\Big)^{2}\simeq 0.919$ via $\text{SNuB}_{\text{corr}}=\text{SNuB}_{\text{Li}}\cdot q\hskip 19.91684pt\text{and}\hskip 19.91684ptL_{B0,\text{corr}}=L_{B0,\text{Li}}\cdot q^{-1}$ in tab. 3. The rates are expressed in units of one SN Ia every 100 years. A plot of the positions of the galaxies with the corresponding rates and distances is shown in fig. 3.
@@
The estimation of cumulative SN Ia rate within 10 Mpc is $31\pm 7$ SN/100 yrs, $245\pm 67$ SN/100 yrs for a 20 Mpc horizon and $650\pm 157$ SN/100 yrs for a 30 Mpc horizon. The errors are computed via a Monte Carlo error estimation starting from the errors on the parameters given by Li et al. (2011), and a 20% correction has been applied to account for the Zone of Avoidance (see Sect. 3.7).
@@
Since the detectable population outside our Galaxy consists only of the most massive DWDs that experience the merging (see Sect. 5), only super-Chandrasekhar merging systems are considered. The realistic number of mergers expected is not sufficient to produce a significant statistics, so a higher number is considered and the results are to be intended as probabilistic, once normalized to the true number of expected events. A broader population is obtained by considering all the merging events produced by the synthetic population in a time span from the nominal observation time $t_{0}=13.5$ Gyr to $t_{0}+2.4$ Myr. The time interval $\Delta t=2.4$ Myr is chosen observing that the merger rate of the population is stable for 2.4 Myr and than drops, due to the fact that the population is depleted from the DWD that merged in the meantime. The merger rate is determined by evolving the separation between the binaries only under GW energy dissipation and considering a Roche scenario as the merging condition. From the resulting population we choose to sample a merging population 500 times larger than the expected one (without incompleteness correction, see Sect. 3.7).
3.7 Error budget